OpenStudy (anonymous):

Find the horizontal asymptote of f(x) = (2 x+5 x^3+3)/(4 x^3-4 x^2-2)

OpenStudy (anonymous):

@Hero @ganeshie8 @mathstudent55 @sammixboo

OpenStudy (anonymous):

I don't know how to approach the problem... but I know what a horizontal asymptote is

OpenStudy (anonymous):

So my initial thought was to look at the degree of the denominator and the numerator but that doesn't seem to work for this one

hero (hero):

To find the horizontal asymptote of a rational function, take the limit as x goes to infinity of the leading terms: \[\lim_{x \rightarrow \infty}\frac{5x^3}{4x^3}\]

OpenStudy (anonymous):

We are not doing limits... that is beyond my skill level

OpenStudy (anonymous):

I am only in precalculus...

OpenStudy (anonymous):

Is there another way to find the horizontal asymptote?

OpenStudy (anonymous):

The textbook says to look at the degree of the numerator and denominator and compare the two... but I'm thinking that either the top or bottom need to be factored

hero (hero):

Watch this video: https://www.youtube.com/watch?v=c-yK2hUnSB0

OpenStudy (anonymous):

OK, I watched that video. His example is an easy one though. I have one that is not as simple as just reducing the leading coefficients of the denominator and numerator

OpenStudy (anonymous):

That is why I posted the problem about 15 hours ago.

hero (hero):

It is that simple if you put the polynomials in standard form

OpenStudy (anonymous):

Rewrite it as 5x^3+2x+3/4x^3-4x^2-2

hero (hero):

Exactly

hero (hero):

\[f(x) = \frac{5x^3 + 2x + 3}{4x^3 - 4x^2 - 2}\]

OpenStudy (anonymous):

The denominator and numerator are both equal so I will take the leading coefficients

OpenStudy (anonymous):

Horizontal asymptote at 5/4

OpenStudy (anonymous):

That was a lot easier than I made it out to be. Thanks, I feel like I should have solved that within 15 hours

OpenStudy (anonymous):

I should have trusted my natural instinct which was to reorder it but that is the same time I posted for help

hero (hero):

The internet is your friend. Use all resources: Google, OS, Khan Academy, Patrick JMT, MIT OCW, Paul's Online Notes, etc.

OpenStudy (anonymous):

Do you have a set of resources that you frequently use? I occasionally use Patrick JMT, Paul's Online Notes, and OS

hero (hero):

I use the whole internet as a resource.